# CoLoRA: Continuous low-rank adaptation for reduced implicit neural modeling of parameterized partial differential equations

### v1

@@ 4 Training CoLoRA models offline
To impose regularity with respect to $t$, we introduce a hyper-network $h:\mathcal{T}\times\mathcal{D}\times\mathbb{R}^{q^{\prime}}\to\mathbb{R}$ that depends on the parameter vector $\boldsymbol{\psi}\in\mathbb{R}^{q^{\prime}}$; see [22, 71, 20] for other methods that build on hyper-networks in different settings. Time $t$ and physics parameter $\boldsymbol{\mu}$ are inputs to the hyper-network and $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ is the output. We focus on the case where $h$ is an MLP, but we stress that any other regression model can be used that provides the necessary regularity from $t$ to $\boldsymbol{\phi}(t,\boldsymbol{\mu})$. For our purposes here, it is sufficient to choose $h$ such that it is continuous in $t$. This means that its output written as $\boldsymbol{\phi}(t;\boldsymbol{\mu})$ will depend continuously on time $t$.
@@
Figure 3: CoLoRA models achieve orders of magnitude lower errors than the empirical best-approximation error of linear model reduction methods, which is in agreement with Section 3.3 that states that CoLoRA parameterizations circumvent the Kolmogorov barrier. Errors are computed on test physics parameters that have not been seen during training.
@@ 5 Online phase of CoLoRA models
Given a new parameter $\boldsymbol{\mu}$ that we have not seen during training, the goal of the online phase is to rapidly approximate the high-fidelity numerical solution $u_{\text{F}}$ at $\boldsymbol{\mu}$. With pre-trained CoLoRA models, we can go about this in two fundamentally different ways: First, we can take a purely data-driven route and simply evaluate the hyper-network $h$ at the new $\boldsymbol{\mu}$ at any time $t$. Second, because the latent state $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ depends on time $t$, we can take an equation-driven route and use the governing equation given in (1) to derive the online parameter $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ via a variational formulation such as Neural Galerkin schemes [15]; see [55, 25, 4, 10] for other sequential-in-time methods that could be combined with CoLoRA.
@@ 5.1 Data-driven forecasting (CoLoRA-D)
We refer to CoLoRA models as CoLoRA-D if predictions at a new physics parameter $\boldsymbol{\mu}$ are obtained by evaluating the hyper-network $h$ at $\boldsymbol{\mu}$ and the times $t$ of interest. The predictions that are obtained from CoLoRA-D models are purely data-driven and therefore do not directly use the governing equations in any way; neither during the pre-training nor during the online phase. Reduced models based on CoLoRA-D are non-intrusive [36, 54], which can have major advantages in terms of implementation and deployment because only data needs to be available; these advantages are the same as for operator learning [61, 62] that also is non-intrusive and typically relies only on data rather than the governing equations. The accuracy of CoLoRA-D models, however, critically depends on the generalization of $h$, which is in agreement with data-driven forecasting in general that has to rely alone on the generalization of data-fitted functions.
@@ 5.2 Equation-driven predictions (CoLoRA-EQ)
If the governing equations given in (1) are available, they can be used together with CoLoRA models to compute the states $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ for a new parameter $\boldsymbol{\mu}$ in a variational sense. We follow Neural Galerkin schemes [15], which provide a method for solving for parameters that enter non-linearly so that the corresponding parameterizations, in a variational sense, solve the given PDE. We stress that such a variational formulation is possible with CoLoRA models because the latent state $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ depends on time $t$ rather than time being an input as the spatial coordinate $\boldsymbol{x}$. In particular, the sequential-in-time training of Neural Galerkin schemes is compatible with the time-dependent online parameter $\boldsymbol{\phi}$. Together with Neural Galerkin schemes, CoLoRA provide solutions that are causal, which is different from many purely data-driven methods.
Neural Galerkin schemes build on the Dirac-Frenkel variational principle [24, 31, 63, 55], which can be interpreted as finding time derivatives $\dot{\boldsymbol{\phi}}(t,\boldsymbol{\mu})$ that solve the Galerkin condition
@@
so that the residual $r_{t}(\boldsymbol{\phi},\dot{\boldsymbol{\phi}})=\partial_{t}{\hat{u}}-f(t,\cdot,{\hat{u}})$ of the PDE (1) as a function over the spatial domain $\Omega$ is orthogonal to the tangent space of the manifold $\{{\hat{u}}(\cdot;\boldsymbol{\theta},\boldsymbol{\phi})\,|\,\boldsymbol{\phi}\in\mathbb{R}^{q}\}$ induced by the online parameters at the current function ${\hat{u}}(\cdot;\boldsymbol{\theta},\boldsymbol{\phi})$; we refer to [15] for details and to Appendix H for the computational procedure.
@@
Figure 4: Left: Purely data-driven CoLoRA (CoLoRA-D) is more than four orders of magnitude faster than traditional numerical models. If the governing equation is solved in a Galerkin-optimal variational sense in the CoLoRA parameterization (CoLoRA-EQ), we still obtain about two orders of magnitude speedups while maintaining causality in the solution. Right: CoLoRA is more data efficient than operator learning and thus well suited for low-data regimes (Burgers’, Vlasov).
@@ 6.1 PDE problems
The following three problems are challenging to reduce with conventional linear model reduction methods because the dynamics are transport dominated [74]. Additional details on these equations and the full order models are provided in Appendix C.
  Collisionless charged particles in electric field The Vlasov equation describes the motion of collisionless charged particles under the influence of an electric field. We consider the setup of [39], which demonstrates filamentation of the distribution function of charged particles as they are affected by the electric field. Our physics parameter $\mu\in[0.2,0.4]$ enters via the initial condition. The full numerical model benchmarked in Figure 4 uses second-order finite differences on a $1024\times 1024$ grid with adaptive time integrator.
@@
  Other PDE models We also look at other PDEs to benchmark against methods from the literature; see Table 1. These include a two-dimensional wave problem with a four-dimensional physics parameter.
@@ 6.2 CoLoRA architectures
The reduced-model parameterization ${\hat{u}}$ is an MLP with CoLoRA layers. The hyper-network $h$ is an MLP with regular linear layers. Both use swish activation functions. The most important architectural choice we make is the size of our networks—${\hat{u}}$ has 8 layers each of width 25 and $h$ has three layers each of width 15. As discussed earlier, such small networks are sufficient because of the strong inductive bias and low-rank continuity in time of CoLoRA networks. The error metric we report is the mean relative error, which is also our loss function (6). All additional choices regarding the architecture, training and data sampling are detailed in Appendix D and Figure 7.
@@ 6.3 CoLoRA and number of latent parameters
First, the linear approximations are ineffective for all three examples, which is in agreement with the observation from Section 6.1 that the three PDE models are challenging to reduce with linear model reduction methods. Second, our CoLoRA models achieve orders of magnitude lower relative errors for the same number of parameters as linear approximations. In all examples, 2–3 latent parameters are sufficient in CoLoRA models, which is in agreement with the low dimensionality of the physics parameters of these models. After the steep drop off of the error until around $q=2$ online parameters, there is a slow improvement if any as we increase $q$, which is in agreement with other nonlinear approximations methods [19, 56]. This is because once $q$ is equal to the intrinsic dimension of the problem, compression no longer helps reduce errors in predictions and instead the error is driven by other error sources such as time integration and generalization of the hyper-network. In these examples, the purely data-driven CoLoRA-D achieves slightly lower relative errors than the equation-driven CoLoRA-EQ, which could be due to the time integration error. In any case, the CoLoRA-EQ results show that we learn representations that are consistent with the PDE dynamics in the Dirac-Frenkel variational sense.
@@ 6.6 Leveraging physics knowledge in the online phase with CoLoRA-EQ
Figure 5: Solving the governing equations in a variational sense with CoLoRA parameterizations (CoLoRA-EQ) leads to causal solutions and allows conserving quantities such as mass in the Vlasov problem, which is key for building trust in physics predictions and for interpretability.
We build on conserving Neural Galerkin schemes introduced in [88] to conserve the mass of the probability distribution that describes the particles in the Vlasov problem. Preserving unit mass can be important for physics interpretations. In Figure 5, we show that using the CoLoRA-EQ with [88], we are able to conserve the mass of solution fields of the Vlasov equation to machine precision. By contrast, neither the CoLoRA-D nor F-FNO conserve the quantity, as the numerical experiments show.
@@ 6.7 Comparison to other nonlinear methods
PDE method rel. err param $q$
2D Wave DINo 6.89e-3 $208,936$ 50
2D Wave CoLoRA-D 9.22e-4 7505 49
1D Burg. CROM 1.93e-3 19470 2
1D Burg. CoLoRA-D 2.74e-04 4521 2
Vlasov F-FNO 8.57e-3 $>10^{6}$
Vlasov CoLoRA-D 9.87e-04 5451 7
2D Burg. F-FNO 5.11e-3 $>10^{6}$
2D Burg. CoLoRA-D 4.96e-04 5451 7
Table 1: CoLoRA is more accurate than a range of other methods for forecasting PDEs and model reduction based on implicit neural representations and operator learning. Notice that $q$ is the latent/online dimension.
We run CoLoRA on the 2D wave benchmark problem described in the DINo publication [96]; see Table 1. CoLoRA achieves an about one order of magnitude lower error than DINo while using about one order of magnitude fewer parameters. We also compare CoLoRA against CROM [19], another nonlinear reduce modeling method that uses neural networks. We benchmark our method against the 1D inviscid Burgers’ problem that the authors of CROM describe in their publication [19]. CoLoRA achieves an about one order of magnitude lower error than CROM, again using about one order of magnitude fewer parameters.
7 Conclusions
CoLoRA leverages that PDE dynamics are typically continuous in time while evolving on low-dimensional manifolds. CoLoRA models provide nonlinear approximations and therefore are efficient in reducing transport-dominated problems that are affected by the Kolmogorov barrier. At the same time, CoLoRA is data efficient and requires only few training trajectories in our examples. The continuous-in-time adaptation of CoLoRA network weights leads to rapid predictions of solution fields of PDEs at new physics parameters, which outperforms current state-of-the-art methods. We provide an implementation of CoLoRA at https://github.com/julesberman/CoLoRA.
@@ Appendix A Literature review of nonlinear model reduction
There is a wide range of literature on model reduction; see [5, 84, 9, 6, 54] for surveys and textbooks. We focus here on model reduction methods that build on nonlinear parameterizations to circumvent the Kolmogorov barrier [74].
First, there is a range of methods that pre-compute a dictionary of basis functions and then subselect from the dictionary in the online phase [46, 23, 2, 26, 65, 75, 47, 33]. However, once the dictionary has been pre-computed offline, it remains fixed and thus such dictionary-based localized model reduction methods are less flexible in this sense compared to the proposed CoLoRA approach.
Second, there are nonlinear reduced modeling methods that build on nonlinear transformations to either recover linear low-rank structure that can be well approximated with linear parameterizations in subspace or that augment linear approximations with nonlinear correction terms. For example, the early work [83] shows how to shift bases to account for translations and other symmetries. Other analytic transformations are considered in, e.g., [69, 81, 27, 77, 72, 8, 34, 45]. The works [91, 16] parameterize the transformation maps and train their parameters on snapshot data rather than using transformations that are analytically available.
Third, there are online adaptive model reduction methods that adapt the basis representation during the online phase [53, 86, 44, 35, 17, 76, 97, 73, 12, 11, 80, 42]. An influential line of work is the one on dynamic low-rank approximations [53, 68, 29, 28, 68, 66, 67, 40] that adapt basis functions with low-rank additive updates over time and thus can be seen as a one-layer version of CoLoRA reduced models.

### v2

@@ 4 Training CoLoRA models offline
To impose regularity with respect to $t$, we introduce a hyper-network $h:\mathcal{T}\times\mathcal{D}\times\mathbb{R}^{q^{\prime}}\to\mathbb{R}^{q}$ that depends on the parameter vector $\boldsymbol{\psi}\in\mathbb{R}^{q^{\prime}}$; see (de Avila Belbute-Peres et al., 2021; Pan et al., 2023; Cho et al., 2023) for other methods that build on hyper-networks in different settings. Time $t$ and physics parameter $\boldsymbol{\mu}$ are inputs to the hyper-network and $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ is the output. We focus on the case where $h$ is an MLP, but we stress that any other regression model can be used that provides the necessary regularity from $t$ to $\boldsymbol{\phi}(t,\boldsymbol{\mu})$. For our purposes here, it is sufficient to choose $h$ such that it is continuous in $t$. This means that its output written as $\boldsymbol{\phi}(t;\boldsymbol{\mu})$ will depend continuously on time $t$.
@@
Figure 3: CoLoRA models achieve orders of magnitude lower errors than the best-approximation error of linear model reduction methods, which is in agreement with Section 3.3 that states that CoLoRA parameterizations circumvent the Kolmogorov barrier.
@@ 5 Online phase of CoLoRA models
Given a new parameter $\boldsymbol{\mu}$ that we have not seen during training, the goal of the online phase is to rapidly approximate the high-fidelity numerical solution $u_{\text{F}}$ at $\boldsymbol{\mu}$. With pre-trained CoLoRA models, we can go about this in two fundamentally different ways: First, we can take a purely data-driven route and simply evaluate the hyper-network $h$ at the new $\boldsymbol{\mu}$ at any time $t$. Second, because the latent state $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ depends on time $t$, we can take an equation-driven route and use the governing equation given in (1) to derive the online parameter $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ via a variational formulation such as Neural Galerkin schemes (Bruna et al., 2024); see (Lasser & Lubich, 2020; Du & Zaki, 2021; Anderson & Farazmand, 2022; Berman & Peherstorfer, 2023) for other sequential-in-time methods that could be combined with CoLoRA.
@@ 5.1 Data-driven forecasting (CoLoRA-D)
We refer to CoLoRA models as CoLoRA-D if predictions at a new physics parameter $\boldsymbol{\mu}$ are obtained by evaluating the hyper-network $h$ at $\boldsymbol{\mu}$ and the times $t$ of interest. The predictions that are obtained from CoLoRA-D models are purely data-driven and therefore do not directly use the governing equations in any way; neither during the pre-training nor during the online phase. Reduced models based on CoLoRA-D are non-intrusive (Ghattas & Willcox, 2021; Kramer et al., 2024), which can have major advantages in terms of implementation and deployment because only data needs to be available; these advantages are the same as for operator learning (Li et al., 2021; Lu et al., 2021) that also is non-intrusive and typically relies only on data rather than the governing equations. The accuracy of CoLoRA-D models, however, critically depends on the generalization of $h$, which is in agreement with data-driven forecasting in general that has to rely on the generalization of data-fitted functions alone.
@@ 5.2 Equation-driven predictions (CoLoRA-EQ)
If the governing equations given in (1) are available, they can be used together with CoLoRA models to compute the states $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ for a new parameter $\boldsymbol{\mu}$ in a variational sense. We follow Neural Galerkin schemes (Bruna et al., 2024), which provide a method for solving for parameters that enter non-linearly so that the corresponding parameterizations, in a variational sense, solve the given PDE. We stress that such a variational formulation is possible with CoLoRA models because the latent state $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ depends on time $t$ rather than time being an input as the spatial coordinate $\boldsymbol{x}$. In particular, the sequential-in-time training of Neural Galerkin schemes is compatible with the time-dependent online parameter $\boldsymbol{\phi}$. Together with Neural Galerkin schemes, CoLoRA provides solutions that are causal, which is different from many purely data-driven methods.
Neural Galerkin schemes build on the Dirac-Frenkel variational principle (Dirac, 1930; Frenkel, 1934; Lubich, 2008; Lasser & Lubich, 2020), which can be interpreted as finding time derivatives $\dot{\boldsymbol{\phi}}(t,\boldsymbol{\mu})$ that solve the Galerkin condition
@@
so that the residual $r_{t}(\boldsymbol{\phi},\dot{\boldsymbol{\phi}})=\partial_{t}{\hat{u}}-f(t,\cdot,{\hat{u}})$ of the PDE (1) as a function over the spatial domain $\Omega$ is orthogonal to the tangent space of the manifold $\{{\hat{u}}(\cdot;\boldsymbol{\theta},\boldsymbol{\phi})\,|\,\boldsymbol{\phi}\in\mathbb{R}^{q}\}$ induced by the online parameters at the current function ${\hat{u}}(\cdot;\boldsymbol{\theta},\boldsymbol{\phi})$; we refer to (Bruna et al., 2024) for details and to Appendix H for the computational procedure.
@@
Figure 4: Left: Purely data-driven CoLoRA (CoLoRA-D) is more than four orders of magnitude faster than traditional numerical models. If the governing equations are solved with Neural Galerkin (Bruna et al., 2024) in a Galerkin-optimal variational sense in the CoLoRA parameterization (CoLoRA-EQ), we still obtain about two orders of magnitude speedups while maintaining causality in the solution. Right: CoLoRA is more data efficient than operator learning and thus well suited for low-data regimes (Burgers’, Vlasov).
@@ 6.1 PDE problems
The following three problems are challenging to reduce with conventional linear model reduction methods because the dynamics are transport dominated (Peherstorfer, 2022). Additional details on these equations and the full order models are provided in Appendix C.
  Collisionless charged particles in electric field The Vlasov equation describes the motion of collisionless charged particles under the influence of an electric field. We consider the setup of Güçlü et al. (2014), which demonstrates filamentation of the distribution function of charged particles as they are affected by the electric field. Our physics parameter $\mu\in[0.2,0.4]$ enters via the initial condition. The full numerical model benchmarked in Figure 4 uses second-order finite differences on a $1024\times 1024$ grid with adaptive time integrator.
@@
  Other PDE models We also look at other PDEs to benchmark against methods from the literature; see Table 1. These include a two-dimensional wave problem with a four-dimensional physics parameter taken from Yin et al. (2023) and a three-dimensional shallow water wave example from Serrano et al. (2023).
@@ 6.2 CoLoRA architectures
The reduced-model parameterization ${\hat{u}}$ is an MLP with CoLoRA layers. The hyper-network $h$ is an MLP with regular linear layers. Both use swish activation functions. The most important architectural choice we make is the size of our networks—${\hat{u}}$ has 8 layers each of width 25 and $h$ has three layers each of width 15. As discussed earlier, such small networks are sufficient because of the strong inductive bias and low-rank continuity in time of CoLoRA networks. Only for the 3D shallow water example we use layer width 128. The larger width helps to capture the oscillations in the solution field in this example; see also Section 7. The error metric we report is the mean relative error, which is also our loss function (6). More details are in Appendix D and Figure 7.
@@ 6.3 CoLoRA and number of latent parameters
First, the linear approximations are ineffective for all three examples, which is in agreement with the observation from Section 6.1 that the three PDE models are challenging to reduce with linear model reduction methods. Second, our CoLoRA models achieve orders of magnitude lower relative errors for the same number of parameters as linear approximations. In all examples, 2–3 latent parameters are sufficient in CoLoRA models, which is in agreement with the low dimensionality of the physics parameters of these models. After the steep drop off of the error until around $q=2$ online parameters, there is a slow improvement if any as we increase $q$, which is in agreement with other nonlinear approximations methods (Chen et al., 2023b; Lee & Carlberg, 2020). This is because once $q$ is equal to the intrinsic dimension of the problem, compression no longer helps reduce errors in predictions and instead the error is driven by other error sources such as time integration and generalization of the hyper-network. In these examples, the purely data-driven CoLoRA-D achieves slightly lower relative errors than the equation-driven CoLoRA-EQ, which could be due to the time integration error. In any case, the CoLoRA-EQ results show that we learn representations that are consistent with the PDE dynamics in the sense of Neural Galerkin based on the Dirac-Frenkel variational principle.
@@ 6.5 Data efficiency versus operator learning
example: three-dim. spherical shallow water two-dim. wave
metric: MSE number parameters MSE number parameters
MP-PDE 9.37e-5 - 9.256e-7 -
DeepONet 6.54e-3 - 1.847e-2 -
DINo 4.48e-5 $2,022,912$ 9.495e-6 $579,776$
CORAL 3.44e-6 $1,049,344$
CoLoRA-D 3.19e-06 $335,744$ 1.891e-07 $7505$
Table 1: CoLoRA is more accurate than a range of other methods for forecasting PDEs and model reduction based on implicit neural representations while using significantly fewer parameters. MSE values and parameter counts are taken from Yin et al. (2023) for the 2D wave problem and from Serrano et al. (2023) for 3D spherical shallow water. Parameter counts are estimated from the descriptions architecture depth and width in the original papers.
@@ 6.6 Leveraging physics knowledge in the online phase with CoLoRA-EQ
Figure 5: Solving the governing equations in a variational sense with Neural Galerkin (Bruna et al., 2024) and CoLoRA parameterizations (CoLoRA-EQ) leads to causal solutions and allows conserving quantities (Schwerdtner et al., 2023) such as mass in the Vlasov problem, which is key for building trust in physics predictions and for interpretability.
We build on conserving Neural Galerkin schemes introduced in Schwerdtner et al. (2023) to conserve the mass of the probability distribution that describes the particles in the Vlasov problem. Preserving unit mass can be important for physics interpretations. In Figure 5, we show that using the CoLoRA-EQ with Schwerdtner et al. (2023), we are able to conserve the mass of solution fields of the Vlasov equation to machine precision. By contrast, neither the CoLoRA-D nor F-FNO conserve the quantity, as the numerical results indicate.
@@ 6.7 Comparison to other nonlinear methods
We run CoLoRA on two benchmark problems. The first is described in the DINo publication (Yin et al., 2023). This is a 2D wave problem where the four dimensional parameter $\mu$ affects the position and magnitude of the initial condition. The second is described in the CORAL publication (Serrano et al., 2023). It is a shallow water equation formulated over a 3D spherical domain where the $\mu$ parameter nonlinearly affects the initial condition. We additionally report the accuracy of two other methods MP-PDE (Brandstetter et al., 2022) and DeepONet (Lu et al., 2021) both originally benchmarked in Yin et al. (2023) and Serrano et al. (2023). We report these results in Table 1. We see that on these two challenging benchmark problems CoLoRA achieves the lowest mean squared error. In terms of implementations, CoLoRA succeeds using a relatively simple modulation scheme and straightforward pre-training. CoLoRA also outperforms all other methods while using close to one to two orders of magnitude fewer parameters.
7 Conclusions, limitations, and future work
CoLoRA leverages that PDE dynamics are typically continuous in time while evolving on low-dimensional manifolds. CoLoRA models provide nonlinear approximations and therefore are efficient in reducing transport-dominated problems that are affected by the Kolmogorov barrier. At the same time, CoLoRA is data efficient and requires only few training trajectories in our examples. The continuous-in-time adaptation of CoLoRA network weights leads to rapid predictions of solution fields of PDEs at new physics parameters, which outperforms current state-of-the-art methods.
Limitations First, the theoretical analysis for reduced models based on CoLoRA is currently limited. The preliminary results on overcoming the Kolmogorov barrier for a specific setup with the linear advection equation cannot be directly generalized to other problems and thus a more in-depth analysis is necessary. Second, there are applications where pre-training the CoLoRA network once and for all is insufficient, such as when predicting bifurcations that are not represented in the training data. Then, an online adaptive updating of the offline parameters is desired, for which efficient methods need to be developed.
Future work First, to well approximate solution fields with high-frequency oscillations, sharp gradients, and other non-smooth features, reduced modeling with CoLoRA can be combined with Fourier feature embeddings and periodic activation functions. Second, our hyper-network based method of modulation succeeds mainly when generalizing to examples which are in-distribution in terms of $\boldsymbol{\mu}$ and $t$. Later work might seek to expand CoLoRA’s method of parameter modulation to settings with neural ordinary differential equations and other methods that can enhance CoLoRA’s extrapolation ability. Third, a future direction is scaling reduced modeling with CoLoRA to higher-dimensional problems in both parameter and spatial domain. We expect that active data collection will be key for CoLoRA models to be efficient in high dimension.
We provide an implementation of CoLoRA at https://github.com/julesberman/CoLoRA.
@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Literature review of nonlinear model reduction
There is a wide range of literature on model reduction; see Antoulas (2005); Rozza et al. (2008); Benner et al. (2015); Antoulas et al. (2021); Kramer et al. (2024) for surveys and textbooks. We focus here on model reduction methods that build on nonlinear parameterizations to circumvent the Kolmogorov barrier (Peherstorfer, 2022).
First, there is a range of methods that pre-compute a dictionary of basis functions and then subselect from the dictionary in the online phase (Jens L. Eftang & Patera, 2011; Dihlmann et al., 2011; Amsallem et al., 2012; Eftang & Stamm, 2012; Maday & Stamm, 2013; Peherstorfer et al., 2014; Kaulmann et al., 2015; Geelen & Willcox, 2022). However, once the dictionary has been pre-computed offline, it remains fixed and thus such dictionary-based localized model reduction methods are less flexible in this sense compared to the proposed CoLoRA approach.
Second, there are nonlinear reduced modeling methods that build on nonlinear transformations to either recover linear low-rank structure that can be well approximated with linear parameterizations in subspace or that augment linear approximations with nonlinear correction terms. For example, the early work Rowley & Marsden (2000) shows how to shift bases to account for translations and other symmetries. Other analytic transformations are considered in, e.g., Ohlberger & Rave (2013); Reiss et al. (2018); Ehrlacher et al. (2020); Qian et al. (2020); Papapicco et al. (2022); Barnett & Farhat (2022); Geelen et al. (2023); Issan & Kramer (2023). The works by Taddei et al. (2015); Cagniart et al. (2019) parameterize the transformation maps and train their parameters on snapshot data rather than using transformations that are analytically available.
Third, there are online adaptive model reduction methods that adapt the basis representation during the online phase (Koch & Lubich, 2007; Sapsis & Lermusiaux, 2009; Iollo & Lombardi, 2014; Gerbeau & Lombardi, 2014; Carlberg, 2015; Peherstorfer & Willcox, 2015; Zahr & Farhat, 2015; Peherstorfer, 2020; Black et al., 2020; Billaud-Friess & Nouy, 2017; Ramezanian et al., 2021; Huang & Duraisamy, 2023). An influential line of work is the one on dynamic low-rank approximations (Koch & Lubich, 2007; Musharbash et al., 2015; Einkemmer & Lubich, 2019; Einkemmer et al., 2021; Musharbash et al., 2015; Musharbash & Nobile, 2017, 2018; Hesthaven et al., 2022) that adapt basis functions with low-rank additive updates over time and thus can be seen as a one-layer version of CoLoRA reduced models.